If one number is half of another and their average is 15, what is the value of the smaller number increased by 2?
(A) 8 (B) 10 (C) 12 (D) 20 (E) 30

Answer by mananth(16946)

let the number be x
the other number = x/2
x+ x/2 / 2 =15
2x+x / 2 /2 =15
3x/4 =15
3x=60
x=20
the other number is 10.
10+2 = 12 is the answer